A 1,200 KG CONCRETE SLAB MEASURES 2m X 2m X 20 cm is delivered to a building under construction. Does it contain steel reinforcing steel bars?


1
Expert's answer
2021-02-17T11:06:31-0500

The destiny is

"\rho=\frac mV=\frac{m}{abc},"

"\rho=\frac{1200}{2\cdot2\cdot 0.2}=1500~\frac{kg}{m^3}" is more than concrete's destiny, so it contains steel bars.